Circular No. 6661

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cn IN NGC 5490 W.-d. Li, Y.-l. Qiu, Q.-y. Qiao, Y. Zhang, W. Zhou, and J.-y. Hu,

Beijing Astronomical Observatory (BAO), on behalf of the BAO Supernova Survey, report the discovery of an apparent supernova (mag ∼ 15.8) on an unfiltered CCD frame taken with the BAO 0.6-m reflector on May 14.6 UT. SN 1997cn is located at α = 14^{h}09^{m}57^{s}.76, δ = +17^{o}32′32′′.3 (equinox 2000.0), which is 6′′.75 east and 11′′.7 south of the nucleus of NGC 5490. The supernova was confirmed on images obtained on May 15. Unfiltered CCD images of the same field taken on Apr. 30 (limiting mag ∼ 18.0) show no star at the position of the supernova.

GRB 970508 M. Mignoli, C. Bartolini, A. Bragaglia, G. Clementini, A. Comastri,

and A. Guarnieri, Bologna Astronomical Observatory and Bologna University; and F. Frontera, L. Nicastro, and E. Costa on behalf of the BeppoSAX team report: “Exposures in the Kron-Cousins R band obtained with the 1.52-m telescope at Loiano on May 10.82–10.88 UT show that Bond’s variable (IAUC 6654) was at R = 19.78 ± 0.05 (rms); no variability is detected among the four exposures (limiting mag R ∼ 21). The observing time corresponds to the maximum suggested by Djorgovski et al. (IAUC 6658), and our measurements provide a better constraint to the lightcurve shape at its peak. Three R exposures obtained on May 9.84–9.89 do not reveal the object (limiting mag R = 20.5). The optical brightening 24 hr after the γ-ray burst is similar to that seen for the GRB 970228 on images obtained at the Loiano telescope. A finding chart from the May 10.85 image is available at

http://www.bo.astro.it.”

M. Garcia, D. Moraru, and J. McClintock, Harvard-Smithsonian Center for Astrophysics; C. R. Robinson and C. Kouveliotou, Universities Space Research Association and Marshall Space Flight Center; and J. van Paradijs, University of Amsterdam and University of Alabama, Huntsville, report: “Observations with the Mount Hopkins 1.2-m telescope in the R band during nonphotometric conditions, spanning May 12.177–12.247 UT, show that the counterpart found by Bond (IAUC 6654) was 0.77 ± 0.09 fainter than the star 13′′ to the north, corresponding to R = 20.47 ± 0.09 using the calibration from Schaefer et al. (IAUC 6658). Eleven consecutive integrations of 300-s duration each are consistent with a steady source having only the expected 0.08-mag statistical fluctuations.”

1997 May 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green